# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v1

@@ 6 Conclusion
In this paper, we introduced a novel methodology for estimating the mediation functional in the presence of high-dimensional covariates and mediators, potentially exceeding the sample size. Our approach allows for the covariate-treatment interaction in the generation of the mediators, as well as both covariate-treatment and mediator-treatment interactions in the generation of the response. The main contribution of our work is a novel debiasing technique in the presence of these two challenges. The structure of the mediation functional, which was our parameter of interest, possesses certain properties which required development of new debiasing methodology. Our proposed methodology can be of independent interest in other statistical settings where the parameter of interest has a structure similar to that of the mediation functional. We showed that our estimator is root-$n$ consistent and asymptotically normal, which aids in performing statistical inference on the direct/indirect treatment effects by providing an asymptotically valid confidence interval. To the best of our knowledge, this is the first study to address the challenge of estimating NDE and NIE when dealing with high-dimensional covariates, mediators, and their interactions with the treatment variable simultaneously. While this paper focuses on a linear model, extending the methodology to generalized linear models is a natural next step, which we leave as an intriguing direction for future research.
@@ Appendix A Proofs
$\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq\|\check{a}-\xi_{1}\|_{\infty}+\|\xi_{2}\|_{\infty}=\max_{1\leq j\leq p+q}\left|\check{a}_{j}-\frac{1}{n_{t}}\sum_{i:A_{i}=1}Q_{i}^{\top}A_{j}Q_{i}\right|+\|\xi_{2}\|_{\infty}\,.$ (A.7)
@@
$\left\|A_{j}\right\|_{F}^{2}=\operatorname{tr}\left(A_{j}^{\top}A_{j}\right)$ $=\operatorname{tr}\left(\Sigma_{t}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{W}^{-\frac{1}{2}}\Sigma_{W}^{-\frac{1}{2}}\check{a}e_{j}^{\top}\Sigma_{W}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(\Sigma_{W}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{W}^{-1}\check{a}e_{j}^{\top}\Sigma_{W}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(e_{j}\Sigma_{W}^{\frac{1}{2}}\Sigma_{t}^{\frac{1}{2}}\check{a}^{\top}\Sigma_{W}^{-1}\check{a}\right)$
$=\left(\Sigma_{W}\right)_{jj}\check{a}^{\top}\Sigma^{-1}_{W}\check{a}\leqslant V_{1}^{\prime}S_{1}\,,$
@@
$\mathbb{P}\left(\left|\frac{1}{n_{t}}\sum_{i:A_{i}=1}Z_{t,i}\right|\geq\sqrt{\frac{\log{(p+q)}}{n_{t}}}\mid\mathcal{F}_{0,n}\right)\leq\frac{\operatorname{Var}\left(Z_{t}\mid\mathcal{F}_{0,n}\right)}{\log{(p+q)}}\,.$
@@
For the first term, we have $\|\bar{X}-\mu\|_{\infty}=K_{6}\sqrt{\log{p}/n_{c}}$ with probability going to 1 for some constant $K_{6}>0$, by a simple application of Hoeffding’s inequality along with a union bound. We next obtain the upper bound of $\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}$ in Equation (A.13). Write $(\mathbf{X}_{c}^{\top}\tau_{2}^{*})$ as:
@@
$\kappa^{2}$ $\leq(\mathbf{W}_{t}^{\top}\tau_{1})_{j}^{2}\leq\|\tau_{1}\|_{2}^{2}\sum_{\{i:A_{i}=1\}}\mathbf{W}_{t,ij}^{2}\implies n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{t}}{n}\frac{1}{n_{t}}\sum_{i:A_{i}=1}\mathbf{W}^{2}_{ij}}\,,$
$\kappa^{2}$ $\leq(\mathbf{X}_{c}^{\top}\tau_{2})_{j}^{2}\leq\|\tau_{2}\|_{2}^{2}\sum_{\{i:A_{i}=0\}}\mathbf{X}_{c,ij}^{2}\implies n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{c}}{n}\frac{1}{n_{c}}\sum_{i:A_{i}=0}\mathbf{X}^{2}_{ij}}\,.$
@@
which follows by cancelling $\sqrt{n}$ from the definition of $\sigma_{n}$. Since the noise is centered i.i.d sub-Gaussian random variables (Assumption 5.5), we have conditional on $\mathcal{G}_{n}=\sigma(\mathcal{D}_{1},\mathbf{X},\mathbf{M}_{c},A_{1:n})$, $T_{1}$ has mean 0 and variance 1. Furthermore, it is easy to check that Lyapunov’s condition is also satisfied with $\delta=1$ as $\sqrt{n}\|\tau_{1}\|_{\infty}+\sqrt{n}\|\tau_{2}\|_{\infty}=o_{p}(1)$ and $\|\gamma_{1}\|_{2}$ is bounded. Therefore, an application of the conditional central limit theorem (Bulinski, (2017)) yields:

### v2

@@ 7 Conclusion
In this paper, we introduced a novel methodology for estimating the mediation functional (and consequently NDE, NIE) in the presence of ultra-high-dimensional covariates and mediators. Our model is sufficiently flexible to accommodate interactions between the treatment assignment and both covariates and mediators in the conditional mean of the response, as well as interactions between the treatment assignment and covariates in the conditional expectation of the mediators. The primary contribution is to develop a novel debiasing framework to obtain a $\sqrt{n}$-CAN estimate of the mediation functional (consequently for NDE, NIE) in the presence of these challenges, which is potentially applicable to other problems sharing a similar structural equation model. To the best of our knowledge, this is the first study to address the challenge of estimating NDE and NIE when dealing with high-dimensional covariates, mediators, and their interactions with the treatment variable simultaneously. While this paper focuses on a linear model, extending the methodology to generalized linear models is a natural next step, which we leave as an intriguing direction for future research.
@@ Appendix A Proofs
$\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq\|\check{a}-\xi_{1}\|_{\infty}+\|\xi_{2}\|_{\infty}=\max_{1\leq j\leq p+q}\left|\check{a}_{j}-\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}A_{j}Q_{i}\right|+\|\xi_{2}\|_{\infty}\,.$ (A.7)
@@
$\left\|A_{j}\right\|_{F}^{2}=\operatorname{tr}\left(A_{j}^{\top}A_{j}\right)$ $=\operatorname{tr}\left(\Sigma_{t}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{t}^{-\frac{1}{2}}\Sigma_{t}^{-\frac{1}{2}}\check{a}e_{j}^{\top}\Sigma_{t}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(\Sigma_{t}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{t}^{-1}\check{a}e_{j}^{\top}\Sigma_{t}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(e_{j}\Sigma_{t}^{\frac{1}{2}}\Sigma_{t}^{\frac{1}{2}}\check{a}^{\top}\Sigma_{t}^{-1}\check{a}\right)$
$=\left(\Sigma_{t}\right)_{jj}\check{a}^{\top}\Sigma^{-1}_{t}\check{a}\leqslant V_{1}^{\prime}S_{1}\,,$
@@
$\mathbb{P}\left(\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Z_{t,i}\right|\geq\sqrt{\frac{\log{(p+q)}}{n_{t}}}\mid\mathcal{F}_{0,n}\right)\leq\frac{\operatorname{Var}\left(Z_{t}\mid\mathcal{F}_{0,n}\right)}{\log{(p+q)}}\,.$
@@
For the first term, we have $\|\bar{X}-\mu\|_{\infty}=K_{6}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1 for some constant $K_{6}>0$, by a simple application of Hoeffding’s inequality along with a union bound. We next obtain the upper bound of $\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}$ in Equation (A.13). Write $(\mathbf{X}_{c}^{\top}\tau_{2}^{*})$ as:
@@
$\kappa^{2}$ $\leq(\mathbf{W}_{t}^{\top}\tau_{1})_{j}^{2}\leq\|\tau_{1}\|_{2}^{2}\sum_{\{i:A_{i}=1\}}\mathbf{W}_{t,ij}^{2}\implies n\|\tau_{1}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{t}}{n}\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}\mathbf{W}^{2}_{ij}}\,,$
$\kappa^{2}$ $\leq(\mathbf{X}_{c}^{\top}\tau_{2})_{j}^{2}\leq\|\tau_{2}\|_{2}^{2}\sum_{\{i:A_{i}=0\}}\mathbf{X}_{c,ij}^{2}\implies n\|\tau_{2}\|_{2}^{2}\geq\frac{\kappa^{2}}{\frac{n_{c}}{n}\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}\mathbf{X}^{2}_{ij}}\,.$
@@
which follows by cancelling $\sqrt{n}$ from the definition of $\sigma_{n}$ in Equation (A.29). Since the noise is centered i.i.d. sub-Gaussian random variables (Assumption 5.5), we have conditional on $\mathcal{G}_{n}=\sigma(\mathcal{D}_{1},\mathbf{X},\mathbf{M}_{c},A_{1:n})$, $T_{1}$ has mean 0 and variance 1. Furthermore, it is easy to check that Lyapunov’s condition is also satisfied with $\delta=1$ as $\sqrt{n}\|\tau_{1}\|_{\infty}+\sqrt{n}\|\tau_{2}\|_{\infty}=o_{p}(1)$ and $\|\gamma_{1}\|_{2}$ is bounded. Therefore, an application of the conditional central limit theorem (Bulinski, (2017)) yields:
@@
In this section, we present details on the calculation of different effects. We will investigate $\mathbb{E}[Y^{(1,M^{(1)})}]$, $\mathbb{E}[Y^{(0,M^{(0)})}]$, $\mathbb{E}[Y^{(1,M^{(0)})}]$, $\mathbb{E}[Y^{(0,M^{(1)})}]$, indirect effect and direct effect.
A.5 Calculation of Effects
By taking the expectation of regression Equations (2.3) for $Y$ and $M$ conditional on $(X,M,A)$ and $(X,A)$ respectively, we obtain the following equations:
$\mathbb{E}[Y^{(1,M^{(1)})}]$ $=\alpha_{1}+\mathbb{E}[X]^{\top}\beta_{1}+\mathbb{E}[M\mid X,A=1]^{\top}\gamma_{1}=(\alpha_{1}+\delta_{1}^{\top}\gamma_{1})+\mathbb{E}[X]^{\top}(\beta_{1}+\mathbf{B}_{1}^{\top}\gamma_{1})\,,$
$\mathbb{E}[Y^{(0,M^{(0)})}]$ $=\alpha_{0}+\mathbb{E}[X]^{\top}\beta_{0}+\mathbb{E}[M\mid X,A=0]^{\top}\gamma_{0}=(\alpha_{0}+\delta_{0}^{\top}\gamma_{0})+\mathbb{E}[X]^{\top}(\beta_{0}+\mathbf{B}_{0}^{\top}\gamma_{0})\,,$
$\mathbb{E}[Y^{(1,M^{(0)})}]$ $=\alpha_{1}+\mathbb{E}[X]^{\top}\beta_{1}+\mathbb{E}[M\mid X,A=0]^{\top}\gamma_{1}=(\alpha_{1}+\delta_{0}^{\top}\gamma_{1})+\mathbb{E}[X]^{\top}(\beta_{1}+\mathbf{B}_{0}^{\top}\gamma_{1})\,,$
$\mathbb{E}[Y^{(0,M^{(1)})}]$ $=\alpha_{0}+\mathbb{E}[X]^{\top}\beta_{0}+\mathbb{E}[M\mid X,A=1]^{\top}\gamma_{0}=(\alpha_{0}+\delta_{1}^{\top}\gamma_{0})+\mathbb{E}[X]^{\top}(\beta_{0}+\mathbf{B}_{1}^{\top}\gamma_{0})\,.$
The indirect effect are as follows:
NIE $=\mathbb{E}[Y^{(1,M^{(1)})}]-\mathbb{E}[Y^{(1,M^{(0)})}]$
$=(\alpha_{1}+\delta_{1}^{\top}\gamma_{1})+\mathbb{E}[X]^{\top}(\beta_{1}+\mathbf{B}_{1}^{\top}\gamma_{1})-(\alpha_{1}+\delta_{0}^{\top}\gamma_{1})-\mathbb{E}[X]^{\top}(\beta_{1}+\mathbf{B}_{0}^{\top}\gamma_{1})$
$=(\delta_{1}-\delta_{0})^{\top}\gamma_{1}+\mathbb{E}[X]^{\top}(\beta_{1}-\beta_{1}+\mathbf{B}_{1}^{\top}\gamma_{1}-\mathbf{B}_{0}^{\top}\gamma_{1})$
$=(\delta_{1}-\delta_{0})^{\top}\gamma_{1}+\mathbb{E}[X]^{\top}(\mathbf{B}_{1}-\mathbf{B}_{0})^{\top}\gamma_{1}\,.$
For direct effect, we have:
NDE $=\mathbb{E}[Y^{(1,M^{(0)})}]-\mathbb{E}[Y^{(0,M^{(0)})}]$
$=(\alpha_{1}+\delta_{0}^{\top}\gamma_{1})+\mathbb{E}[X]^{\top}(\beta_{1}+\mathbf{B}_{0}^{\top}\gamma_{1})-(\alpha_{0}+\delta_{0}^{\top}\gamma_{0})-\mathbb{E}[X]^{\top}(\beta_{0}+\mathbf{B}_{0}^{\top}\gamma_{0})$
$=(\alpha_{1}-\alpha_{0}+\delta_{0}^{\top}(\gamma_{1}-\gamma_{0}))+\mathbb{E}[X]^{\top}(\beta_{1}+\mathbf{B}_{0}^{\top}\gamma_{1}-\beta_{0}-\mathbf{B}_{0}^{\top}\gamma_{0})$
$=\alpha_{1}-\alpha_{0}+\delta_{0}^{\top}(\gamma_{1}-\gamma_{0})+\mathbb{E}[X]^{\top}(\beta_{1}-\beta_{0})+\mathbb{E}[X]^{\top}\mathbf{B}_{0}^{\top}(\gamma_{1}-\gamma_{0})\,.$
Appendix B Sensitivity Analysis
$k_{1}=k_{2}=5$ RMSE
$p+q$ $K=1$ $K=1.25$ $K=1.5$ $K=1.75$ $K=2$ $K=2.25$ $K=2.5$ $K=2.75$ $K=3$
$100$ 0.2754 0.2606 0.2782 0.2631 0.2269 0.2255 0.2230 0.2065 0.2143
$800$ 0.1779 0.1534 0.1464 0.1427 0.1330 0.1237 0.1261 0.1167 0.1224
$1500$ 0.2747 0.2600 0.2731 0.2546 0.2531 0.1978 0.1976 0.1660 0.1763
$2500$ 0.1425 0.1385 0.1214 0.1123 0.1117 0.1101 0.1102 0.1062 0.1153
Table 4: Sensitivity Analysis of $(K_{1},K_{2})$ in RMSE over 400 simulation replications $n=1250$ and $\sigma^{2}=0.5$ (The three lowest RMSE in each row have been boldfaced).
$k_{1}=k_{2}=5$ RMSE
$p+q$ $K=1$ $K=1.25$ $K=1.5$ $K=1.75$ $K=2$ $K=2.25$ $K=2.5$ $K=2.75$ $K=3$
$100$ 0.1340 0.1288 0.1200 0.1153 0.1119 0.1074 0.1085 0.1115 0.1122
$800$ 0.1784 0.1500 0.1441 0.1380 0.1283 0.1319 0.1270 0.1279 0.1281
$1500$ 0.2203 0.2041 0.1747 0.1696 0.1498 0.1575 0.1472 0.1460 0.1465
$2500$ 0.2752 0.2664 0.2812 0.2796 0.3054 0.2274 0.2570 0.2344 0.2646
Table 5: Sensitivity Analysis of $(K_{1},K_{2})$ in RMSE over 400 simulation replications $n=1000$ and $\sigma^{2}=0.5$ (The three lowest RMSE in each row have been boldfaced).
$k_{1}=k_{2}=5$ RMSE
$p+q$ $K=1$ $K=1.25$ $K=1.5$ $K=1.75$ $K=2$ $K=2.25$ $K=2.5$ $K=2.75$ $K=3$
$100$ 0.2644 0.2442 0.2121 0.2035 0.1910 0.1936 0.1822 0.1773 0.1535
$800$ 0.3308 0.3157 0.2811 0.2511 0.2573 0.2915 0.2925 0.2795 0.2953
$1500$ 0.4311 0.3213 0.2375 0.2048 0.2004 0.1672 0.1827 0.1719 0.1722
$2500$ 0.2312 0.2147 0.2171 0.2279 0.3119 0.3282 0.3135 0.2499 0.2754
Table 6: Sensitivity Analysis of $(K_{1},K_{2})$ in RMSE over 400 simulation replications $n=750$ and $\sigma^{2}=0.5$ (The three lowest RMSE in each row have been boldfaced).
$n=1250$ RMSE
$p+q=1500$ $K_{1}=1$ $K_{1}=1.25$ $K_{1}=1.5$ $K_{1}=1.75$ $K_{1}=2$ $K_{1}=2.25$ $K_{1}=2.5$ $K_{1}=2.75$ $K_{1}=3$
$K_{2}=1$ 0.3957 0.3815 0.3929 0.3679 0.3919 0.3852 0.3849 0.3729 0.3640
$K_{2}=1.25$ 0.3985 0.3867 0.3653 0.3953 0.3941 0.3833 0.3708 0.3959 0.3891
$K_{2}=1.5$ 0.3170 0.3288 0.3065 0.3241 0.3020 0.3144 0.3082 0.3058 0.3242
$K_{2}=1.75$ 0.2368 0.2272 0.2410 0.2315 0.2283 0.2500 0.2293 0.2439 0.2417
$K_{2}=2$ 0.2094 0.2084 0.2055 0.2101 0.2214 0.2297 0.2250 0.2266 0.2266
$K_{2}=2.25$ 0.1983 0.1935 0.2018 0.2063 0.1989 0.2120 0.1954 0.1888 0.2103
$K_{2}=2.5$ 0.1862 0.1818 0.1852 0.2027 0.1951 0.1975 0.1988 0.2154 0.2017
$K_{2}=2.75$ 0.1848 0.1768 0.1884 0.1955 0.1808 0.1878 0.1832 0.1722 0.1808
$K_{2}=3$ 0.1697 0.1700 0.1690 0.1762 0.1762 0.1713 0.1833 0.1843 0.1722
Table 7: Sensitivity Analysis of different $(K_{1},K_{2})$ in RMSE over 400 simulation replications $n=1250$, $p+q=1500$ and $\sigma^{2}=0.5$ (The five lowest RMSE have been boldfaced).